Prev Next

SW&IT Meeting Minutes 2020-06-15

Attendees: Carl, David, Igor, Matthieu, Stephen, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20200601
LR20200601 - Bec, Matthieu 80.0
LR20200601 - Holmberg, Carl 56.0
LR20200601 - Edmeades, David 72.0
LR20200601 - Sola, Igor 80.0
LR20200601 - Hooper, Stephen 80.0
LR20200601 - Zhang, Yang 80.0
LR20200601 - Zhang, Xianyu 80.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

MB

  • Last week
    • OC, Management
    • TMS/sci-arch meetings
    • review nightwriter/peepshow with Shane
    • review lbttools iraf conversion to Python3 with Al, Alexander and Michelle
    • LN/Argos/SOUL/PB discussion with Al
    • SHARK-VIS finished RFA
    • Student interviews and selection
    • more AnyDesk tests with Olga, ROWG related discussion emails/slack
  • This week
    • Applicant interviews
    • DMS
      • websockets, AdSec, etc.
      • LBC stiched png previewfigure
      • max_bytes for ovms
      • merge IMS into DMS
    • cobbler / kickstart, ims f30
    • RPM/CI short tutorial
    • TMS CoDR GMT
    • ESO Esspresso review

DE

  • Cisco training
  • Macbook setup--ready to go. Timing for Carl?
  • PH email list changes
  • PIT machine httpd logs
    • Quick manual fix for operations
    • Work on method (logrotate/cron/whatever) to automate
    • Mountian machine as well--at least zip them, but ideally push to Tucson (rsync?)
  • Luci hardware document
  • CDW followup
  • New laptop for ML; trying once, Quixote style, to convince Property that $200 for an 8 year old laptop is more than $1. Will give up soon and order
  • JC/AO shared drive discussion died? I'm ready to help if they need it
  • Anydesk operation help to get observers ready for first night of observing
Next, continue with training, documents, sensor cals, etc.

YZ

  • Previous:
    • GCSGUI new patch (v2.1), partially fixed its problems.
    • TMS (general) meeting, plus some TMS wiki work.
    • Some AO-Cravat work (wiki/cleanup/code).
  • Next:
    • Continue GCSGUI investigation (now a new IT#8124).
    • TMS software development (higher priority, needs to be ready for closed-dome test sometime later).
    • AO-Cravat work continue (wiki/cleanup/code), meeting on Friday morning.
    • Other work if possible (MODS, Python etc.).
    • Safety training if time permits.

CH

  • Last Week:
    • AHS: LBTI alarm patch
    • Argos: continue work on RPM build.
    • TCS, etc: update leap second file, per IT7319.
    • Safety training
  • This Week:
    • Argos: continue work on RPM build.
    • MODS: check if there's a more elegant way to check MODS AGW and guide camera connectivity after restarts.
    • AHS: LBTI alarm patch, incorporate updated INDI code, or use more sophisticated text processing.

SH

  • yesterweek:
    • video card and DisplayPort MST hub for testing
    • Particle monitor.
    • zabbix for Olga and MODS.
    • accounts for most of the non-LBTO personell -- work on cutover for VPN -- going to be some pain.
    • TCS stuff with Josh.
    • obs1 stuff with Jay -- reboot 06/14/2020
    • remote observing: x2go, anydesk. Discussions and x2go issues from rm507.
    • luci x2go -- bashrc changed
    • add development packages to shell for Christian.
  • next:
    • Second monitor for rm538 (hopefully)
    • Fedora burn on machine capable of subbing for four-headed monster -- try it with the Matrox card (maybe it will work).
    • Satellite images, weather serial into TCS (arrange testing), particle monitor telemetry (and how)
    • Fedora 32 RPM builds -- commited, but I have asked that alternatives be sought (I have expressed my reservations about Fedora -- and CentOS).

IS

Last week:
  • LUCI
    • I installed the software also in robs.mountain.lbto.org for Fix #IT7952
    • LUCI WebIO device http API (ongoing)
  • PIT
    • PIT admin tool backend work, mostly DevObs work at this point. I want to have the whole full stack development in a docker container. Easier for other people to develop on it as they don't have to install anything in their local dev machines. Also to distribute it to production when it's ready.
This week:
  • Safety Training
  • LUCI
    • Partner with Barry/Dave on testing #IT7952?
    • work on configuration of LUCI WebIO, the HTTP GET api seems to be the same but on this device seems to be not consistent. Have to work a bit more on it. David E. provided support for network installation and backup configuration if needed.
    • Need to partner with Jenny and revert/test to old GEIRS in LUCI2 osa-luci.sim
  • PIT
    • Work on the PIT admin tool backend.
  • OT
    • Partner with Andrew to work on MODS XSLT

XZ

  • Last week
    • AO/SOUL operation support
    • new 4D operation support
    • SX AdSec issues
    • Power Backplane action items discussion
  • This week
    • AO operation support?
    • new 4D support?
    • SX AdSec issues
    • Power Backplane action items
Topic revision: r12 - 15 Jun 2020, XianyuZhang
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback